OpenStudy (anonymous):

improper integral∫_0^∞▒〖e^x/(e^2x+3) dx〗

myininaya (myininaya):

First we need to evaluate: \[\int\limits_{}^{}\frac{e^x}{e^{2x}+3} dx\] \[\int\limits_{}^{}\frac{e^x}{(e^x)^2+3} dx\] You need to recall a few trig identities and see which one will be useful here: \[(\sin(\theta))^2=1-(\cos(\theta))^2\] \[(\sec(\theta))^2=1+(\tan(\theta))^2\] \[(\tan(\theta))^2=(\sec(\theta))^2-1\]

myininaya (myininaya):

both terms in the bottom have + in front of it so we will use the 2nd trig identity I mentioned, right?

myininaya (myininaya):

\[(\sec(\theta))^2=1+(\tan(\theta))^2 => (a \sec(\theta))^2=a^2+(a \tan(\theta))^2=(a \tan(\theta))^2+a^2\] So we have: \[(a \sec(\theta))^2=(a \tan(\theta))^2+a^2\] So what do you think our substitution would look like?

myininaya (myininaya):

What is a?

myininaya (myininaya):

\[\text{ and what is } \tan(\theta) ?\]

OpenStudy (anonymous):

lol

myininaya (myininaya):

We have that we are trying to make the bottom \[(a \tan(\theta))^2+a^2 \] so we can write it as that one term \[(a \sec(\theta))^2\] So we are trying to figure out the substitution such that \[(e^x)^2+3=(a \tan(\theta))^2+a^2\]

myininaya (myininaya):

So what is the substitution that we need here \[e^x=?\]

OpenStudy (mertsj):

\[e^x=a \tan \theta\]

myininaya (myininaya):

Gj mertsj

myininaya (myininaya):

now what is a?

OpenStudy (anonymous):

\[u=e^x\] \[du=e^xdx\] \[\int \frac{du}{u^2+3}\] almost done

OpenStudy (mertsj):

ty. I thought someone should reward you for all that hard work.

myininaya (myininaya):

And satellite likes to remember extra things

OpenStudy (mertsj):

\[a=\sqrt{3}\]

myininaya (myininaya):

So he is going to show you a "short cut"

myininaya (myininaya):

But my way shows why his way is true

OpenStudy (anonymous):

now look in te back of the book to find integrals of the form \[\frac{du}{u^2+a^2}\] and you will see that it is \[\frac{1}{a}\tan^{-1}(\frac{u}{a})\]

OpenStudy (mertsj):

I think Satellite is a "get ur dun" guy

OpenStudy (anonymous):

because all this nonsense is just showing off. look i can integrate this, look i can integrate that... the entire content of the majority of calc 2 is on the back two pages of the text

OpenStudy (mertsj):

You and I, on the other hand, are ladies of refinement and class.

OpenStudy (anonymous):

refinement, class, and reinvent the wheel

OpenStudy (mertsj):

That too.